black and white are codominant in a group of kittens. the group consists of 20 black kittens, 24 gray kittens, and 50 white kittens. let B represent the allele for black and b represent the allele for white.
- what are genotypes and allele frequencies.
- is population in H-W Equilibrium?
Answer #1

Here BB = Black = 20

Bb = Grey = 24

bb = whitw = 50

Total kittens = 94

1. For genotype frequency we use formula : Individual of specific genotype/ Total genotype:

for BB = 20/94 = 0.22

for Bb= 24/94 = 0.25

for bb = 50/94 = 0.53

1. For allele frequency we use formual : total Individuals of specfic allele/ Total allele in population:

Here BB = 20 = 40 allele of B

Bb = 24 = 24 allele of B and 24 allele of b

bb = 50 = 100 allele of b

For B allel frequency = total Individuals of B allele/ Total allele in population = 64/188 = 0.34

For b allel frequency = total Individuals of b allele/ Total allele in population = 124/188 = 0.66

2. A population in genetic equilibrium after a generation of mating will have the following genotype frequencies:

p2     2pq     q2

BB Bb bb.

To find if our population is in H-W equilibrium we will compare observed and expected genotype frequencies.

For expected.

p2 = BB = 0.34 * 0.34 = 0.12

2pq = Bb = 2 * 0.34 * 0.66 = 0.44

q2 = bb = 0.66 * 0.66 = 0.43

Our observed and expected genotype frequency are different, so popluation is not in H-W equilibrium.